Approximately 6.3 million wrecked cars are recycled in two consecutive years. There were 100,000 more recycled the second year than the year How many cars are recycled each year ?
Let x be the number of cars, in millions, recycled in the first year. Write an equation for the total numbers of cars recycled.

Answer by ikleyn(52785)

.

x + (x+0.1) = 6.3


2x = 6.3 - 0.1

2x = 6.2

 x = 6.2/2 = 3.1.


ANSWER.  3.1 million cars were recycled in 1st year. 

         3.2 million cars were recycled in 2nd year.

Solved.